mary wants to make 200 pounds of blended coffee selling for $5.50 per pound. How many pounds of $6.00 coffee must she mix with how many pounds of $4.00?
Answer by ankor@dixie-net.com(22740)

mary wants to make 200 pounds of blended coffee selling for $5.50 per pound.
How many pounds of $6.00 coffee must she mix with how many pounds of $4.00?
:
Let x = amt of $6 coffee required
the total is given as 200 lb, therefore:
(200-x) = amt of $4 coffee
:
A typical mixture equation:
6x + 4(200-x) = 5.50(200)
:
:
You should be able to do it now. If not, let me know in the comments.